# FEUDA: Frustratingly Easy Prompt Based Unsupervised Domain Adaptation

### v1

@@ Appendix G More On Analyzing the Impact of MLM on UDA
Table 12: Impact of Masking Rate on FEUDA. We train FEUDA using varying random masking rates on the Apparel $\to$ Movies domain pair. With high masking rates, the performance on the source domain is largely maintained, but the performance on the target domain rapidly deteriorates.
Appendix H Learning in a Few-Shot Setup
Classification Accuracy
Table 13 accompanies Figure 3 (Section 6.3), showing the 256-shot performance of FEUDA and other baselines, across model sizes.
Model Src Only Src+Tgt MMD Two Phase MMD FEUDA
Source Target Source Target Source Target Source Target Source Target
T5v1.1 Base 92.0 (0.4) 79.3 (0.8) 93.2 (0.1) 82.8 (0.6) 80.1 (0.7) 62.5 (0.7) 92.3 (0.5) 79.8 (1.4) 90.1 (0.5) 81.2 (0.7)
T5v1.1 XL 91.1 (6.1) 80.9 (8.4) 95.7 (0.1) 92.5 (0.4) 87.3 (6.8) 71.7 (7.8) 92.2 (0.1) 84.3 (0.9) 95.2 (0.3) 86.8 (2.2)
T0 3B 95.3 (0.3) 91.3 (0.3) 95.6 (0.3) 91.8 (0.6) 91.3 (5.1) 79.5 (6.7) 54.0 (4.8) 53.5 (0.4) 95.8 (0.1) 92.8 (0.2)
Table 13: Performance of FEUDA across different models, in a 256-shot learning setup on the Apparel $\to$ Movies domain pair. We see FEUDA retaining strong performance on the target domain across models.
Impact of Masked Words
We extend the analysis on the impact of masked word selection from Section 6.2 to the few-shot setting in Table 14, where we compare the impact of masking informative or uninformative words. We also consider two different few-shot setups: one with access to the full unlabelled datasets in phase 1 pre-training, and another where even the unlabelled data is few-shot. Similar to the full-data setting, random masking remains most powerful, indicating that both semantic and background features are necessary for effective classification on the unlabelled domain. However, unlike the full-data setting where informative and uninformative masking are comparable, in the few-shot setting, informative masking is significantly more useful.
@@ Appendix H Learning in a Few-Shot Setup
Table 14: Impact of word selection for masking, in a 256-shot learning setup. We evaluate FEUDA on the Apparel $\to$ Movies domain pair, and select words for masking based on their “informativeness” to the classification task. Random masking is most powerful for the target domain, indicating that both semantic and background features are necessary for effective classification on the unlabelled domain. However, informative masking is significantly more useful than uninformative masking.
Appendix I Comparison With More Baselines
Our main comparisons are made with the MMD based method proposed by Malik et al. (2023), as it is a recent and powerful UDA method that outperforms other popular invariance based UDA approaches like DANN Ganin et al. (2016); Li et al. (2018) and DSN Bousmalis et al. (2016).
[…]
Additionally, with an emerging class of weight interpolation based methods, we make a comparison with task vector arithmetic Ilharco et al. (2022). The use of task vectors with PEFT has been unexplored in the literature, and we find that the method does not work with IA3. With fully fine-tuned models, the method improves in performance, but is still weaker than FEUDA.
@@ Appendix I Comparison With More Baselines
FEUDA 93.7 (0.3) 83.3 (0.9)
MMD 94.7 (0.3) 81.3 (0.6)
DANN 53.5 (2.3) 52.3 (1.7)
CORAL 94.6 (0.2) 80.9 (0.4)
Wasserstein 94.5 (0.2) 82.5 (0.4)
Task Vectors 46.3 (0.3) 48.0 (0.7)
Task Vectors (fine-tuning) 93.0 (0.2) 69.0 (0.4)
[…]
Appendix J More About MMD
The Maximum Mean Discrepancy (MMD) Gretton et al. (2012) measures the difference between first order moments of variables in a Reproducing Kernel Hilbert Space Aronszajn (1950). Multiple lines of work have shown that minimizing divergence measures like MMD, when combined with auxiliary task-specific loss functions, results in training instabilities and vanishing gradients Kashyap et al. (2021); Han and Eisenstein (2019).
@@
We compare variants of the MMD method in Table 16 and show that the loss is sensitive to small changes in the loss design. Specifically we compare the MMD method used in the main paper with:
@@
FEUDA remains more powerful than all variants.
@@
FEUDA 93.7 (0.3) 83.3 (0.9)
MMD 94.7 (0.3) 81.3 (0.6)
MMD over Logits 95.0 (0.2) 81.3 (0.7)
@@
Two Phase MMD 90.1 (0.1) 68.7 (2.0)
Table 16: Comparison of variants of minimizing MMD, on the Apparel $\to$ Movies domain pair. FEUDA remains more powerful than all variants.
Figure 5: UMap visualizations of sentence embeddings from the Apparel $\to$ Movies data pair, using the T5v1.1 base model and (IA)${}^{3}$ PEFT method. Training with MMD risks stability issues, and all embeddings from the target domain can be mapped to the closest source class cluster. This results in poor classification performance on the target domain.

### v2

@@ Appendix G Impact of Target Domain Exposure
Table 12: Impact of Masking Rate on CPT. We train CPT using varying random masking rates on the Apparel $\to$ Movies domain pair. With high masking rates, the performance on the source domain is largely maintained, but the performance on the target domain rapidly deteriorates.
Appendix H Understanding how CPT aids UDA
Table 13 (accompanies Figure 7) shows the impact of masking sequences at inference, on classification accuracy. Words are selected for masking based on their their “informativeness”, measured by their PMI to the inference class label. The performance of the model is best with the original unmasked sequences, indicating the presence of both informative and uninformative words are essential for strong classification performance.
Table 15 accompanies Figure 7 and shows the impact of varying masking strategies on classification performance, in a few-shot setting. We also consider two different few-shot setups: one with access to the full unlabelled datasets in phase 1 pre-training, and another where even the unlabelled data is few-shot.
To isolate any effects of PEFT methods or pre-training data, we repeat the analysis from Table 15 in Table 14 with fine-tuning Flan-T5 in a full data setting, and note similar trends.
Method Accuracy
Source Target
Original 93.5 83.3
Informative Masking 88.0 78.8
Uninformative Masking 92.0 79.0
Table 13: Impact of masking at inference. We evaluate CPT on the Apparel $\to$ Movies domain pair, and select words for masking based on their “informativeness” to the classification task.
Masking Strategy Accuracy
Source Target
Random 95.8 (0.0) 86.8 (0.3)
Informative 93.9 (0.6) 85.3 (0.3)
Uninformative 95.0 (0.0) 84.8 (0.1)
Table 14: Impact of word selection for masking during training, using Flan-T5 base and no PEFT methods.
@@ Appendix H Understanding how CPT aids UDA
Table 15: Impact of word selection for masking, in a 256-shot learning setup. We evaluate CPT on the Apparel $\to$ Movies domain pair, and select words for masking based on their “informativeness” to the classification task. Random masking is most powerful for the target domain, indicating that both semantic and background features are necessary for effective classification on the unlabelled domain. However, informative masking is significantly more useful than uninformative masking.
Appendix I Single Phase CPT Training
Our proposed approach in Section 2 involves two stages of training, which is more expensive than standard single phase UDA approaches. In this section, we propose a single training phase variant to CPT, and show that it performs similarly to the original method. We use the two phase pipeline in our experiments in the main paper, but note that the single and two phase pipelines are interchangeable.
$(\lambda\;l(\mathbb{C}(x,y);\theta)$
$+(1-\lambda)\;l(\mathbb{M}(x^{\prime});\theta))$
Table 16 compares the performance of the single phase and two phase variants of CPT. We also compare with a vanilla joint single phase objective, where $\lambda$ is fixed at 0.5 through training (called Single Phase Vanilla). The performance of the single and two phase variants are almost identical, and either can be used interchangeably. In comparison, the vanilla single phase method is significantly weaker on the target domain.
@@ Appendix I Single Phase CPT Training
Two Phase 93.7 (0.3) 83.3 (0.9)
Singe Phase 93.5 (0.4) 83.3 (0.5)
Singe Phase Vanilla 93.6 (0.1) 75.0 (5.7)
Table 16: Comparison of single and two-phase variants of CPT, on the Apparel $\to$ Movies domain pair. The single and two phase variants are almost identical in performance.
Appendix J Instability of Domain Invariance Methods for UDA
The Maximum Mean Discrepancy (MMD) Gretton et al. (2012) measures the difference between first order moments of variables in a Reproducing Kernel Hilbert Space Aronszajn (1950). Multiple lines of work have shown that minimizing divergence measures like MMD, when combined with auxiliary task-specific loss functions, results in training instabilities and vanishing gradients Kashyap et al. (2021); Han & Eisenstein (2019).
Figure 9: UMap visualizations of sentence embeddings from the Apparel $\to$ Movies data pair, using the T5v1.1 base model and (IA)3 PEFT method. Training with UDAPTER risks stability issues, and all embeddings from the target domain can be mapped to the closest source class cluster. This results in poor classification performance on the target domain.
@@
We compare variants of the UDAPTER method in Table 17 and show that the loss is sensitive to small changes in the loss design. Specifically we compare the UDAPTER method used in the main paper with:
@@
CPT remains more powerful than all variants.
@@
CPT 93.7 (0.3) 83.3 (0.9)
UDAPTER 94.7 (0.3) 81.3 (0.6)
UDAPTER over Logits 95.0 (0.2) 81.3 (0.7)
@@
Two Phase UDAPTER 90.1 (0.1) 68.7 (2.0)
Table 17: Comparison of variants of minimizing MMD, on the Apparel $\to$ Movies domain pair. CPT remains more powerful than all variants.